# Distributional Instruments: Identification and Estimation with Quantile Least Squares

### v1

@@ 1 Introduction
Our first contribution is to show that distributional relevance is a meaningful notion of instrument strength, not just a descriptive property. Within a nonseparable triangular model, we characterize a class of purely distributional instruments: variables $Z$ for which $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ but $F_{X\mid Z}(\cdot\mid Z)$ is non-degenerate in $L^{2}$. Building on the control-function framework of Chernozhukov et al. (2020), we prove that such instruments still identify average structural effects via the control variable $V=F_{X}(X\mid Z)$, and we relate this notion to existing concepts of instrument strength based on heteroskedasticity (Lewbel, 1997). Thus, even when policy variables are “purely distributional” for a risk index, they can support point identification of its causal effect on outcomes.
Our second, and main, contribution is to develop a practical IV estimator tailored to these designs, which we call Quantile Least Squares (Q–LS). Instead of summarizing the first stage with a conditional mean, Q–LS builds instruments from the entire conditional quantile process of $X$ given $Z$. Concretely, we: (i) estimate a series of first-stage quantile regressions of $X$ on $Z$; (ii) form a data-driven linear combination of these conditional quantiles that best predicts $X$ in mean square error; and (iii) use this optimal quantile-aggregated prediction as the instrument in an otherwise standard linear IV estimator. Under distributional relevance and a mild non-degeneracy condition, the resulting optimal Q–LS instrument is always relevant in the classical sense, even when the conditional mean of $X$ is flat in $Z$. In a “best-case” world where $X\mid Z$ is a pure location shift, we show that Q–LS collapses to optimal 2SLS and attains the same efficiency bound. In more general cases, Q–LS remains reliable precisely when mean-based instruments become weak.
@@ 2.1 Weak identification in triangular and linear IV models
where $Z$ are excluded instruments, $W$ are included exogenous regressors, and $\mathbb{E}[u\mid Z,W]\neq 0$ while $\mathbb{E}[v\mid Z,W]=0$. Classical weak-instrument problems arise when the first-stage coefficients $\pi$ are small, so that the concentration parameter is close to zero and the instruments explain only a tiny fraction of the variation in $X$; see Staiger and Stock (1997), Stock and Yogo (2005), and Dufour (1997) for canonical analyses. In this case, 2SLS and related estimators exhibit large finite-sample bias toward the OLS estimand, heavy-tailed or multimodal sampling distributions, and nonstandard asymptotic behavior ((Andrews and Mikusheva, 2014; Andrews et al., 2019) for a general overview). These issues are amplified in many-instrument settings (e.g. Bekker, 1994; Hansen et al., 2008), where the effective first-stage signal per instrument becomes small even if the joint $F$ is sizable.
@@
We therefore do not claim that Q–LS is generally robust in the weak-IV sense of Staiger and Stock (1997) or that it uniformly dominates Anderson–Rubin or conditional tests (e.g. Andrews et al., 2019; Olea and Pflueger, 2013) when instruments are arbitrarily weak. Rather, the contribution of our distributional relevance framework is to highlight a different margin of strength: designs in which $\operatorname{Var}(\mathbb{E}[X\mid Z])$ is small or zero, so that instruments appear weak for the mean, may still be strong for the distribution, with $\mathbb{E}[h_{\mathrm{opt}}(Z)^{2}]$ bounded away from zero. In such cases, Q–LS leverages distributional variation in $X\mid Z$ to construct an effective instrument $h(Z)$ that satisfies a strong-IV condition even when the original mean-based instrument fails conventional first-stage diagnostics.
Our notion of distributional relevance can be viewed as specifying a “strong-instrument” regime for triangular models at the distributional level: it requires that $F_{X\mid Z}(\cdot\mid z)$ varies nontrivially with $Z$ in an $L^{2}$ sense, even when the conditional mean $\mathbb{E}[X\mid Z]$ is flat. This allows us to focus on designs where instruments may be weak for linear first stages but remain informative about the distribution of $X$, and hence can still support identification and efficient estimation once the first stage is recast in distributional terms.
@@ 4.5.2 Q–LS versus 2SLS in a location model
General comparison
Outside of the pure location case, Q–LS remains consistent under distributional relevance even when mean relevance fails or the conditional mean is misspecified, because it exploits shifts in the entire distribution of $X\mid Z$. In settings where both mean relevance and distributional relevance hold and the efficient IV $h^{*}(Z)$ lies in the closure of the quantile-generated class $\mathcal{H}$, Q–LS again attains the same efficiency bound as optimal 2SLS. In more general cases where $h^{*}(Z)\notin\mathcal{H}$, Q–LS is efficient within $\mathcal{H}$ but may be slightly less efficient than an oracle 2SLS that can use arbitrary functions of $Z$ as instruments. The main gain of Q–LS is robustness: it can remain strongly identified in settings where the classical first-stage mean shift is close to zero but the distribution of $X\mid Z$ varies substantially with $Z$.
@@ 6 Simulation study
The following parts of the data generating processes (DGPs) are the same in all three designs;
@@
Table 2 presents the results for Design A and Design C for sample sizes of 500 and 1000.555Full results for Design A and Design C are available upon request. It is important to note that classical IV is correctly specified, as $Z_{2}$ is strongly mean-relevant in both Design A and Design C. For Design A, we find that all Q-LS variants perform similarly to classical IV, regardless of the number of quantiles used. This is due to the quantile projection being similar to the mean projection. The analytical Q-LS variants provide better coverage (closer to 0.95) than the OLS plug-in standard errors, which tend to yield over-coverage, indicating conservative standard errors in this design. The control function version (DR-CF) performs worse in terms of bias, RMSE, and coverage than all Q-LS variants and classical IV, regardless of sample size. However, its performance does improve as the number of quantiles increases in this design.
@@ A.5 Semiparametric estimation with distributionally relevant instruments
Appendix B Detailed Proofs of Lemma and Propositions

### v2

@@ 1 Introduction
Our first contribution is to show that distributional relevance is a meaningful notion of instrument strength, not just a descriptive property. Within a non-separable triangular model, we characterize a class of purely distributional instruments: variables $Z$ for which $\operatorname{Var}(\mathbb{E}[X\mid Z])=0$ but $F_{X\mid Z}(\cdot\mid Z)$ is non-degenerate in $L^{2}$. Building on the control-function framework of Chernozhukov et al. (2020), we prove that such instruments still identify average structural effects via the control variable $V=F_{X}(X\mid Z)$, and we relate this notion to existing concepts of instrument strength based on heteroskedasticity (Lewbel, 1997). Thus, even when policy variables are “purely distributional” for a risk index, they can support point identification of its causal effect on outcomes.
Our second, and main, contribution is to develop a practical IV estimator tailored to these designs, which we call Quantile Least Squares (Q–LS). [^4] Our notion of “distributional instruments” is orthogonal to the shift-share (Bartik) literature (see (Adao et al., 2019; Borusyak et al., 2025; Goldsmith-Pinkham et al., 2020) for a detailed review). There, the focus is on constructing a composite instrument from many shocks and exposure shares. Here, we take the instrument $Z$ as given and show how to optimally exploit its distributional impact on $X$ via conditional quantiles. In particular, Q–LS does not impose a shift–share structure on $Z$; it can be applied whether $Z$ is a simple policy indicator or a more complex shock. Instead of summarizing the first stage with a conditional mean, Q–LS builds instruments from the entire conditional quantile process of $X$ given $Z$. Concretely, we: (i) estimate a series of first-stage quantile regressions of $X$ on $Z$; (ii) form a data-driven linear combination of these conditional quantiles that best predicts $X$ in mean square error; and (iii) use this optimal quantile-aggregated prediction as the instrument in an otherwise standard linear IV estimator. Under distributional relevance and a mild non-degeneracy condition, the resulting optimal Q–LS instrument is always relevant in the classical sense, even when the conditional mean of $X$ is flat in $Z$. In a “best-case” world where $X\mid Z$ is a pure location shift, we show that Q–LS collapses to optimal 2SLS and attains the same efficiency bound. In more general cases, Q–LS remains reliable precisely when mean-based instruments become weak.
@@ 2.1 Weak identification in triangular and linear IV models
Closest in spirit to our work is the recent “Distributional Instrumental Variable (DIV)” method of Holovchak et al. (2025). They also exploit distributional shifts in $X\mid Z$ and show how to identify and estimate the entire interventional distribution of the outcome using flexible generative models in a nonlinear IV setting. DIV provides conditions under which the interventional distribution is identified, and illustrates designs where 2SLS fails but distributional information suffices for identification and accurate estimation of mean and quantile treatment effects. Relative to this approach, we focus on average structural effects in a triangular model and on the construction of an optimal scalar instrument for linear IV estimands. Our Q–LS procedure can be viewed as a low-dimensional, analytically tractable counterpart to DIV: it concentrates the distributional information in a finite-dimensional quantile dictionary and delivers an optimal $L^{2}$-projection that can be used within the classical IV toolkit. Moreover, when the endogenous regressor is (or can be reduced to) a binary treatment, the Q–LS index provides a scalar monotone instrument, so that the resulting 2SLS coefficient admits a standard LATE interpretation as a positive-weighted average of complier effects across index increments.
@@
where $Z$ are excluded instruments, $W$ are included exogenous regressors, and $\mathbb{E}[u\mid Z,W]\neq 0$ while $\mathbb{E}[v\mid Z,W]=0$. Classical weak-instrument problems arise when the first-stage coefficients $\pi$ are small, so that the concentration parameter is close to zero and the instruments explain only a tiny fraction of the variation in $X$; see Staiger and Stock (1997), Stock and Yogo (2005), and Dufour (1997) for canonical analyses. In this case, 2SLS and related estimators exhibit large finite-sample bias toward the OLS estimand, heavy-tailed or multimodal sampling distributions, and nonstandard asymptotic behavior (see Andrews and Mikusheva, 2014; Andrews et al., 2019, for a general overview). These issues are amplified in many-instrument settings (e.g. Bekker, 1994; Hansen et al., 2008), where the effective first-stage signal per instrument becomes small even if the joint $F$-statistic is sizable.
@@
We therefore do not claim that Q–LS is generally robust in the weak-IV sense of Staiger and Stock (1997) or that it uniformly dominates Anderson–Rubin or conditional tests (e.g. Andrews et al., 2019; Olea and Pflueger, 2013) when instruments are arbitrarily weak. Rather, the contribution of our distributional relevance framework is to highlight a different margin of strength: designs in which $\operatorname{Var}(\mathbb{E}[X\mid Z])$ is small or zero, so that instruments appear weak for the mean, may still be strong for the distribution, with $\mathbb{E}[h_{\mathrm{opt}}(Z)^{2}]$ bounded away from zero. In such cases, Q–LS leverages distributional variation in $X\mid Z$ to construct an effective instrument $h(Z)$ that satisfies a strong-IV condition even when the original mean-based instrument fails conventional first-stage diagnostics. Our notion of distributional relevance thus specifies a “strong-instrument” regime at the distributional level and complements both the generative DIV approach and the classical weak-IV toolkit. This allows us to focus on designs where instruments may be weak for linear first stages but remain informative about the distribution of $X$, and hence can still support identification and efficient estimation once the first stage is recast in distributional terms.
@@ 4.5.2 Q–LS versus 2SLS in a location model
General comparison and LATE interpretation
Outside of the pure location case, Q–LS remains consistent under distributional relevance even when mean relevance fails or the conditional mean is misspecified, because it exploits shifts in the entire distribution of $X\mid Z$. In designs where both mean relevance and distributional relevance hold and the efficient IV $h^{*}(Z)$ lies in the closure of the quantile-generated class $\mathcal{H}$, Q–LS again attains the same efficiency bound as optimal 2SLS. When $h^{*}(Z)\notin\mathcal{H}$, Q–LS is efficient within $\mathcal{H}$ but may be slightly less efficient than an oracle 2SLS that can use arbitrary functions of $Z$ as instruments. The main gain of Q–LS is robustness: it can remain strongly identified in settings where the classical first-stage mean shift is close to zero but the distribution of $X\mid Z$ varies substantially with $Z$.
When the endogenous regressor is (or can be reduced to) a binary treatment $X=D\in\{0,1\}$, Q–LS also admits a familiar LATE interpretation under additional structure. In the positive-normalized version, the Q–LS index $H=h(Z)$ is a convex combination of conditional quantiles of $D\mid Z$, which collapses possibly high-dimensional instruments into a single scalar encouragement variable. Under standard IV validity and a no-defiers monotonicity condition with respect to $H$, the 2SLS coefficient using $H$ as the sole instrument can be written as a positive-weighted average of local average treatment effects for units whose treatment status changes as $H$ increases. Thus, in binary-treatment applications, Q–LS can be viewed not only as a strength/efficiency device but also as a way to construct a scalar monotone instrument that delivers a conventional LATE interpretation (see Appendix B for details).
In the binary-treatment case, the scalar Q–LS index also admits a LATE representation with nonnegative weights on adjacent “index-complier” groups (Appendix B), and these weights can be estimated in practice, so the researcher can diagnose which parts of the instrument support contribute most to the Q–LS effect. In addition, because the Q–LS index is built as a convex combination of conditional quantiles, the estimated weight vector $\hat{w}$ reveals which parts of the $D\mid Z$ distribution (e.g. central vs. tail quantiles) contribute most to instrument strength.
@@ 5.2 Sparse Q–LS via LASSO
5.3 Positive normalized Q–LS weights
Our baseline Q–LS instrument is a linear functional of the conditional quantile process,
$h_{\omega}(Z)=\int_{0}^{1}\omega(\tau)\,Q_{X\mid Z}(\tau\mid Z)\,d\tau,$
with $\omega\in L^{2}(0,1)$. In finite samples, we approximate this object using a quantile grid $0<\tau_{1}<\cdots<\tau_{K}<1$ and fitted quantiles $\hat{g}_{k}(Z)\approx\widehat{Q}_{X\mid Z}(\tau_{k}\mid Z)$, so that $h(Z)\approx\hat{G}(Z)^{\prime}w$ for $\hat{G}(Z)=(\hat{g}_{1}(Z),\ldots,\hat{g}_{K}(Z))^{\prime}$.
For interpretability and stability, we also consider a constrained version that enforces positive, normalized weights:
$w_{\mathrm{PN}}\in\arg\min_{w\in\mathbb{R}^{K}}\ \mathbb{E}\!\left[(X-\hat{G}(Z)^{\prime}w)^{2}\right]\quad\text{s.t.}\quad w_{k}\geq 0\ \ \forall k,\qquad\mathbf{1}^{\prime}w=1.$ (22)
Equivalently, in population form, this corresponds to restricting $\omega(\tau)\geq 0$ and $\int_{0}^{1}\omega(\tau)\,d\tau=1$.
Interpretation.
Under these constraints, $h_{\mathrm{PN}}(Z)=\hat{G}(Z)^{\prime}w_{\mathrm{PN}}$ is an L-statistic: it is a convex combination of conditional quantiles, i.e. a “quantile average” index. Positivity also preserves distributional ordering: if the conditional distribution of $X$ under $Z=z_{2}$ first-order stochastically dominates that under $Z=z_{1}$ (so that $Q_{X\mid Z=z_{2}}(\tau)\geq Q_{X\mid Z=z_{1}}(\tau)$ for all $\tau$), then $h_{\mathrm{PN}}(z_{2})\geq h_{\mathrm{PN}}(z_{1})$ for any admissible weight function. This rules out sign-cancellation across quantiles and yields an index whose movements track distributional shifts in a single direction.
Finite-sample stability.
Because extreme quantiles are noisier to estimate, unconstrained Q–LS may assign large positive and negative weights that amplify estimation error. The simplex constraint $w\geq 0$, $\mathbf{1}^{\prime}w=1$ regularizes the instrument, bounding its variability and often improving robustness when $\hat{G}$ is generated (e.g. via quantile regression). In our empirical implementation we compute $w_{\mathrm{PN}}$ by quadratic programming, and we can combine it with cross-fitting to further limit overfitting of $\hat{G}$.
@@ 6 Simulation study
The following parts of the data-generating processes (DGPs) are the same in all three designs;
@@
Table 2 presents the results for Design A and Design C for sample sizes of 500 and 1000.666Full results for Design A and Design C are available upon request. It is important to note that classical IV is correctly specified, as $Z_{2}$ is strongly mean-relevant in both Design A and Design C. For Design A, we find that all Q-LS variants perform similarly to classical IV, regardless of the number of quantiles used. This is due to the quantile projection being similar to the mean projection. The analytical Q-LS variants provide better coverage (closer to 0.95) than the OLS plug-in standard errors, which tend to yield over-coverage, indicating conservative standard errors in this design. The control function version (DR-CF) performs worse in terms of bias, RMSE, and coverage than all Q-LS variants and classical IV, regardless of sample size. However, its performance does improve as the number of quantiles increases in this design.
@@ A.5 Semiparametric estimation with distributionally relevant instruments
Appendix B A LATE interpretation for Q–LS under a binary treatment
This appendix provides a potential-outcome interpretation of Q–LS as a LATE estimator under additional structure. The main text treats Q–LS primarily as a strength/efficiency device for IV. Here we show that when the endogenous variable is binary (or can be interpreted as a binary treatment derived from a latent index), Q–LS naturally delivers a scalar monotone instrument for which the standard LATE logic applies.
B.1 Setup: binary treatment, scalar instrument, and a Q–LS index
Let $D\in\{0,1\}$ be the endogenous treatment and let $Y(d)$ denote potential outcomes for $d\in\{0,1\}$. We observe $Y=Y(D)$. For the LATE interpretation in this appendix, we specialize to a scalar instrument $Z$ (which can be binary, $Z\in\{0,1\}$, or more generally ordered). If exogenous covariates $Z_{1}$ are present, we work after partialling them out and suppress $Z_{1}$ in the notation.
In the quantile-dictionary implementation of Q–LS, we consider a finite grid $\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}\subset(0,1)$ and define
$G(Z):=\big(Q_{D\mid Z}(\tau_{1}\mid Z),\ldots,Q_{D\mid Z}(\tau_{K}\mid Z)\big)^{\prime},$
where $Q_{D\mid Z}(\tau\mid Z)$ denotes the conditional $\tau$-quantile of $D$ given $Z$. Q–LS then forms scalar indices of the form
$H_{w}(Z):=G(Z)^{\prime}w,\qquad w\in\Delta_{K}:=\big\{w\in\mathbb{R}^{K}:\ w_{k}\geq 0,\ \mathbf{1}^{\prime}w=1\big\}.$
When $w\in\Delta_{K}$, $H_{w}(Z)$ is a convex combination of conditional quantiles—an L-statistic—which is useful both for monotonicity and for interpretability.
Assumption A.4 (IV validity).
(Independence and exclusion) The instrument $Z$ is independent of the potential outcomes and potential treatments, and affects $Y$ only through $D$:
$\{Y(0),Y(1),D(z)\}_{z}\;\perp\!\!\!\perp\;Z.$
(If covariates $Z_{1}$ are included, the independence is conditional on $Z_{1}$.)
Assumption A.5 (Quantile relevance).
(Distributional relevance of the scalar instrument) There exists at least one $\tau_{k}\in\mathcal{T}$ such that the conditional quantile $Q_{D\mid Z}(\tau_{k}\mid Z)$ is not almost surely constant in $Z$. Equivalently, not all components of $G(Z)$ are almost surely flat in $Z$.
Remark 1.
When $D$ is binary, the conditional quantiles $Q_{D\mid Z}(\tau\mid Z)$ are step functions of the propensity score $p(Z):=\Pr(D=1\mid Z)$. In this case, Assumption A.5 is equivalent to $p(Z)$ not being almost surely constant. We formulate relevance in terms of conditional quantiles to keep the connection with the distributional-instrument framework used in the main text.
Quantile relevance ensures that there exist weights $w\in\Delta_{K}$ for which $\mathrm{Cov}\big(H_{w}(Z),D\big)\neq 0$, i.e. the scalar index $H_{w}$ is a relevant instrument for $D$.
Assumption A.6 (Quantile monotonicity in the scalar instrument).
For each $\tau\in\mathcal{T}$, the conditional quantile $z\mapsto Q_{D\mid Z}(\tau\mid z)$ is weakly increasing in $z$.
Lemma A.1 (Convex combinations of quantiles preserve monotonicity).
Under Assumption A.6, any index $H_{w}(Z)=G(Z)^{\prime}w$ with $w\in\Delta_{K}$ is weakly increasing in the scalar instrument $Z$.
Proof.
For each $\tau_{k}\in\mathcal{T}$, $z\mapsto Q_{D\mid Z}(\tau_{k}\mid z)$ is weakly increasing by Assumption A.6. A convex combination of weakly increasing functions is weakly increasing, so $z\mapsto H_{w}(z)=\sum_{k=1}^{K}w_{k}Q_{D\mid Z}(\tau_{k}\mid z)$ is also weakly increasing. ∎
In what follows, we fix a weight vector $w^{\star}\in\Delta_{K}$ (e.g. the one selected by Q–LS) that satisfies Assumption A.5, and write $H:=H_{w^{\star}}(Z)$.
Assumption A.7 (Monotone treatment selection w.r.t. the index).
(No defiers in the index ordering) There exists a version of the potential treatment process $\{D(h)\}_{h}$ such that
$h_{1}\geq h_{0}\quad\Longrightarrow\quad D(h_{1})\geq D(h_{0})\qquad\text{a.s.}$
In particular, if $H_{1}\geq H_{0}$ then $D(H_{1})\geq D(H_{0})$ almost surely.
Assumption A.7 is the usual no-defiers (monotonicity) condition, expressed with respect to the scalar Q–LS index $H$. Assumption A.6 and Lemma A.1 justify interpreting $H$ as a one-dimensional “encouragement" variable inherited from the scalar instrument $Z$.
B.2 Wald LATE using the Q–LS index
When $H$ takes at least two values $h_{0}<h_{1}$ with positive probability, the standard Wald ratio based on those two values identifies a LATE for compliers defined relative to $H$.
Proposition A.1 (LATE for two index values).
Suppose Assumptions A.4, A.5, A.6, and A.7 hold. Let $H$ take the values $h_{0}<h_{1}$ in its support with $\Pr(H=h_{j})>0$ for $j\in\{0,1\}$ and $\mathbb{E}[D\mid H=h_{1}]\neq\mathbb{E}[D\mid H=h_{0}]$. Then
$\frac{\mathbb{E}[Y\mid H=h_{1}]-\mathbb{E}[Y\mid H=h_{0}]}{\mathbb{E}[D\mid H=h_{1}]-\mathbb{E}[D\mid H=h_{0}]}=\mathbb{E}\!\left[Y(1)-Y(0)\ \big|\ D(h_{1})>D(h_{0})\right].$ (38)
Proof. Under IV validity, $\mathbb{E}[Y\mid H=h]=\mathbb{E}[Y(D(h))]$ and $\mathbb{E}[D\mid H=h]=\mathbb{E}[D(h)]$. Write $Y(D(h))=Y(0)+D(h)\{Y(1)-Y(0)\}$ and take differences between $h_{1}$ and $h_{0}$. Monotonicity implies $D(h_{1})-D(h_{0})\in\{0,1\}$, which delivers (38). ∎
B.3 2SLS with a multi-valued scalar index: positive LATE weights
In practice, Q–LS produces a scalar index $H$ with many support points, and empirical work typically uses 2SLS with $H$ (possibly along with controls). While multiple instruments can generate non-convex (possibly negative) weights across different LATEs, a key advantage of using a single scalar index is that the IV estimand can be expressed as a positive-weighted average of LATEs across index increments.
Assume $H$ has finite support $\{h_{1}<\cdots<h_{J}\}$ for exposition, and define adjacent-index LATEs
$\text{LATE}_{j}:=\mathbb{E}\!\left[Y(1)-Y(0)\ \big|\ D(h_{j+1})>D(h_{j})\right],\qquad j=1,\ldots,J-1.$
Let $\Delta_{j}:=\mathbb{E}[D\mid H=h_{j+1}]-\mathbb{E}[D\mid H=h_{j}]\geq 0$ under Assumption A.7.
Proposition A.2 (2SLS as a positive average of adjacent LATEs).
Suppose Assumptions A.4, A.5, A.6, and A.7 hold and $H$ takes values $\{h_{1}<\cdots<h_{J}\}$ with positive probability. Then the population IV coefficient using the single instrument $H$,
$\beta_{IV}(H):=\frac{\mathrm{Cov}(H,Y)}{\mathrm{Cov}(H,D)},$
can be written as
$\beta_{IV}(H)=\sum_{j=1}^{J-1}\delta_{j}\,\text{LATE}_{j},\qquad\delta_{j}:=\frac{\pi_{j}}{\sum_{\ell=1}^{J-1}\pi_{\ell}},\qquad\pi_{j}\geq 0,$
where $\text{LATE}_{j}$ is the adjacent-index LATE between $h_{j}$ and $h_{j+1}$. In particular, the weights $\delta_{j}$ are nonnegative and sum to one.
Remark. The exact form of $\pi_{j}$ depends on the distribution of $H$ (and on whether $H$ is centered), but the key point is conceptual: with a single scalar index satisfying monotonicity, the IV estimand aggregates local LATEs using positive weights tied to how much the index shifts treatment take-up across adjacent levels.
B.4 Why Q–LS is helpful for a LATE interpretation
Propositions A.1A.2 show that the LATE interpretation hinges on (i) validity of the underlying instrument and (ii) monotonicity with respect to a scalar index. Q–LS contributes by constructing such a scalar index from distributional features of the first stage.
(i) Scalar reduction.
Even when the original instrument vector is high-dimensional, the quantile dictionary $G(Z)$ and the Q–LS weights $w^{\star}$ collapse information into a one-dimensional index $H=H_{w^{\star}}(Z)$. This facilitates monotonicity assumptions and avoids the negative-weight issues that can arise when combining multiple instruments in 2SLS.
(ii) Positive normalized weights and monotone distributional shifts.
When $w\in\Delta_{K}$, $H_{w}(Z)$ is a convex combination of conditional quantiles. Under Assumption A.6, higher values of the scalar instrument $Z$ shift each conditional quantile of $D$ weakly to the right, so $H_{w}(Z)$ inherits this ordering (Lemma A.1). This provides a transparent route for arguing that $H$ acts as a monotone “encouragement” variable, supporting Assumption A.7.
(iii) Strength.
Within the class of admissible indices (e.g. $w\in\Delta_{K}$), Q–LS chooses weights to maximize predictive content for the first stage. Thus, Q–LS does not change the causal content of IV validity, but can materially improve finite-sample performance by increasing $\mathrm{Cov}(H,D)$.
Remark 2 (Inspecting which complier groups matter most).
The representation in Proposition A.2 is not only conceptual: in empirical work the weights $\omega_{j}$ can be estimated from the data, because they are proportional to the contribution of each adjacent index increment to $Cov(H,D)$. Concretely, letting $\hat{\Delta}_{j}:=\mathbb{E}_{n}[D\mid H=h_{j+1}]-\mathbb{E}_{n}[D\mid H=h_{j}]$ and $\hat{p}_{j}:=\Pr_{n}(H\in\{h_{j},h_{j+1}\})$, we can form empirical analogues $\hat{\pi}_{j}$ and $\hat{\delta}_{j}$ and report them as a diagnostic. This makes it possible to see which “complier bands” between $h_{j}$ and $h_{j+1}$ contribute most to the overall Q–LS estimate.
Remark 3 (Which parts of the distribution drive the Q–LS index?).
Because the Q–LS index takes the form $H_{w}(Z)=G(Z)^{\prime}w$ with $G(Z)=\big(Q_{X\mid Z}(\tau_{1}\mid Z),\ldots,Q_{X\mid Z}(\tau_{K}\mid Z)\big)^{\prime}$ and $w\in\Delta_{K}$, the estimated weights $\hat{w}$ provide a direct diagnostic of which quantile components are most influential in the first stage. Large weights on upper quantiles (e.g. $\tau\geq 0.8$) indicate that Q–LS is mainly leveraging tail risk, whereas mass on intermediate or lower quantiles reflects sensitivity of the effect to central or lower parts of the $X\mid Z$ distribution. In applications, reporting $\hat{w}$ (or aggregations such as the total weight above a given quantile) helps interpret whether the Q–LS instrument is primarily a location, dispersion, or tail-risk device.
B.5 Scope and limitations
The LATE interpretation in this appendix requires a binary treatment (or a binary treatment derived from a latent/continuous exposure) and monotonicity with respect to the scalar index $H$. When the endogenous regressor is genuinely continuous, the analogous interpretation is in terms of weighted averages of local causal responses (e.g. “marginal treatment effects” or “local IV” objects), rather than LATE. In that case, Q–LS still operates as an instrument-construction/strength device, but the exact causal interpretation is no longer a complier LATE.
Appendix C Detailed Proofs of Lemma and Propositions